The one-sample t statistic for a test of

H0: μ = 10

Ha: μ < 10


Based on n = 10 observations has the value t = -2.25.

a. What are the degrees of freedom for this statistic?

b. What is the P-value for this test?
(4 points)
Answer #1

this is the two tailed test .  

The null and alternative hypothesis is ,

H0 :   = 10

Ha : < 10

df=n - 1 =10-1=9

Test statistic = t =-2.25

P(t< t-2.25 ,9) = 0.0256

P-value = 0.0256